40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2048     Entry time: Mon Oct 5 02:51:08 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: almost there 

Working well tonight: the handoff of CARM to RF (REFL2I), successful reduction of CARM offset to zero, and transition control of MCL path to the OUT1 from the common mode board.  All that's left in lock acquisition is to try and get the common mode bandwidth up and the boost on.

ELOG V3.1.3-